Pencils costing 5 cents each are mixed with 4 cent pencils to form a single pack to sale at $1.04. How many of each type should be used if the number at 4 cents is twice of the number at 5 cents.
Answer by richwmiller(17219)

5x+4y=104
y=2x
x=8 5 cent pencils
y=16 4 cent pencils
check
5*8+4*16=104
40+64=104
104=104
ok